OpenStudy (anonymous):

Does someone know how to solve this problem? T=2(3.14)√1.63/32 Show work please.

OpenStudy (anonymous):

i just gave you the answer....

OpenStudy (anonymous):

factor the term 32 = 4*8 = 4*4*2 and take it out from the root

OpenStudy (anonymous):

Now I understand lol!

OpenStudy (anonymous):

so just ask, don't repost the same quaestion, ....)

OpenStudy (anonymous):

So far I got T=25.12√1.63/2 what do I do now?

OpenStudy (anonymous):

you can devide 1.63/2 =0,815 and then just use calculator to get the square root

OpenStudy (anonymous):

Do I then multiply that with 25.12?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh ohkay now I understand, thank you so much :D

OpenStudy (anonymous):

you welcome